OpenStudy (anonymous):

Will fan and give a medal!! T=1/7 (x-y+z) i need help solving for X

OpenStudy (seascorpion1):

\[T=\frac{x}{7}-\frac{y}{7}+\frac{z}{7}\]\[\frac{x}{7}=T+\frac{y}{7}-\frac{z}{7}\]therefore: \[x=7T+y-z\]